OpenStudy (anonymous):

Multiply the quantity (8 X 10^4)X (21 X 10^-2). Express the answer in scientific notation.

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

1.0x10^-16 1.0x10^16 10x10^15 10x10^16 These are The answer choices

OpenStudy (anonymous):

first you would make it \[8000 \times \frac{ 21 }{ 100 }\]

OpenStudy (anonymous):

then it would be \[21\times 80\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

1680

OpenStudy (anonymous):

\[(8*10^{4})*(21*10^{-2})\]

OpenStudy (anonymous):

is that what it looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait it should be 16800

OpenStudy (anonymous):

i got 16800

OpenStudy (anonymous):

I did 21×80 and got 1680

OpenStudy (anonymous):

its 80000*.21

OpenStudy (anonymous):

Oh okay now i got it but then how do I put that in scientific notation @kola908

OpenStudy (anonymous):

\[10^{1}=10\]\[10^{2}=100\]\[10^{3}=1000\]\[10^{4}=10000\]

OpenStudy (anonymous):

oh okay so it should be 168 x 10^2

OpenStudy (anonymous):

\[16800=1.6*10^{4}\]

OpenStudy (anonymous):

oooh okay

OpenStudy (anonymous):

So it would be B

OpenStudy (anonymous):

1.68X10^4 or 168X10^2

OpenStudy (anonymous):

is there more to the question?

OpenStudy (anonymous):

Okay so it is B because I put down 168x10^2 and got it wrong thanks for all of your help

OpenStudy (anonymous):

And no that is all there is

OpenStudy (phi):

8 X 10^4)X (21 X 10^-2 multiply the numbers out front 8*21 to get 168 multiply the 10^4 * 10^-2 which is easy: add the exponents: 4+ -2 = 2 you get 168 * 10^2 now move the decimal point of 168 two places to the left. compensate by increasing the exponent by 2

OpenStudy (anonymous):

Okay That helps alot Thanks to both of you!!!